# ECoDepth: Effective Conditioning of Diffusion Models for Monocular Depth Estimation

### v1

@@ 3.2 Our Architecture
Table 4: Effectiveness of different embeddings for guiding the diffusion process for depth estimation on NYU Depth v2 dataset. In the third row, rather than using pseudo caption embeddings generated from the scene label (as implemented by VPD [54]), we provide the one-hot vector representing the scene label as a condition to the diffusion model. We observe a slight improvement in the metrics.
@@ 4.3 Ablation Study
Effect of Contextual Information.
As highlighted in the motivation, a key observation of this study is the richness of information contained within the output probability vector of the ViT, surpassing the textual embeddings employed in current state-of-the-art (e.g., VPD [54]). To compare with the utilization of scene label information (as implemented by VPD), we construct the conditioning embedding as a one-hot vector using the scene label, and subsequently transform it using an MLP. As illustrated in Tab. 4, we observe a slight improvement over VPD, suggesting that pseudo-captions are less effective than explicitly providing the class label. Furthermore, our proposed method surpasses both approaches.
@@ 5 Conclusion
We presented a new architecture block Comprehensive Image Detail Embedding (CIDE) module for robust monocular depth estimation in this paper. Our key idea is to do away with intermediate text generation for use with foundational model. Instead, we aim to use richer class-wise probability generated by a classification model, such as ViT. The hypothesis is that, while textual embedding typically highlight salient objects, class-wise probability vector preserves more details, including smaller objects in the background also. We implement the idea using proposed module cascaded with a conditional diffusion pipeline for monocular depth estimation. We demonstrated the effectiveness of our approach on several benchmark datasets and showed that it outperforms SOTA methods by a significant margin.
Acknowledgement. We thank Kartik Anand for assistance with the experiments. This work has been supported by Staqu Technologies Pvt. Ltd., and Department of Science and Technology (DST), Government of India.

### v3

@@ 3.2 Our Architecture
Table 4: Effectiveness of different embeddings for guiding the diffusion process for depth estimation on NYU Depth v2 dataset. In the third row, rather than using pseudo caption embeddings generated from the scene label (as implemented by VPD [54]), we provide the one-hot vector representing the scene label as a condition to the diffusion model. We observe a slight improvement in the metrics, highlighting that information content in the caption is similar to that in one-hot label.
@@ 4.3 Ablation Study
Effect of Contextual Information. As highlighted in the motivation, a key observation of this study is the richness of information contained within the output probability vector of the ViT, surpassing the textual embeddings employed in current state-of-the-art (e.g., VPD [54]). To compare with the utilization of scene label information (as implemented by VPD), we construct the conditioning embedding as a one-hot vector using the scene label, and subsequently transform it using an MLP. As illustrated in Tab. 4, we observe a slight improvement over VPD, indicating that the information content in pseudo-captions resembles that of one-hot labels. Furthermore, our proposed method surpasses both the approaches.
@@ 5 Conclusion
We presented a new architecture block Comprehensive Image Detail Embedding (CIDE) module for robust monocular depth estimation in this paper. Our key idea is to highlight the limitations of using pseudo-captions to provide contextual information, and instead propose to use richer class-wise probability generated by a classification model, such as ViT. The motivation is that, while textual embedding typically highlight salient objects, class-wise probability vector preserves more details, including smaller objects in the background also. We implement the idea using proposed CIDE module cascaded with a conditional diffusion pipeline for monocular depth estimation. We demonstrated the effectiveness of our approach on several benchmark datasets and showed that it outperforms SOTA methods by a significant margin.
Acknowledgement. We thank Kartik Anand for assistance with some of the experiments. This work has been supported by funding through Staqu Technologies Pvt. Ltd., and Department of Science and Technology (DST), Government of India. The travel for Suraj Patni is supported through Research Acceleration funds of Department of Computer Science and Engineering, IIT Delhi.